# Phase diagram and coarsening dynamics in dry apolar active nematics with reciprocal local alignment

### v2

@@ Abstract
Using the Lebwohl-Lasher interaction for reciprocal local alignment, we present a comprehensive phase diagram for a dry, apolar, active nematic system using its stochastic dynamics. The nematic-isotropic transition in this system is first-order and fluctuation-dominated. Our phase diagram identifies three distinct regions based on activity and orientational noise relative to alignment strength: a homogeneous isotropic phase, a nematic phase with giant density fluctuations, and a coexistence region. Using mean-field analysis and hydrodynamic theory, we demonstrate that reciprocal interactions lead to a density fluctuation-induced first-order transition and derive a phase boundary consistent with numerical results. Quenching from the isotropic to nematic phase reveals coarsening dynamics with nematic ordering preceding particle clustering and similar scaling behaviors for density and nematic fields, exhibiting dynamic exponents between 2 and 3.
@@ 1 Introduction
Active nematics represent a fascinating class of nonequilibrium systems characterized by the collective motion and organization of self-propelled entities [1, 2, 3, 4, 5, 6]. Inspired by a myriad of biological systems, ranging from the dynamic organization of the cytoskeleton within cells to the emergent behaviors observed in tissues and even granular materials, active nematics have emerged as a pivotal area of study at the interface of physics, biology, and materials science [4, 7, 3, 8, 9]. At the heart of active systems lies the concept of self-propulsion, where individual constituents continually convert internal energy into directed motion. This intrinsic activity, prevalent in biological systems such as motile cells and bacterial colonies, fuels a rich array of dynamic phenomena, including the spontaneous formation of orientational order, collective motion, and the emergence of complex spatiotemporal patterns [4, 7, 8]. Moreover, active nematics find analogs in nonliving systems, including vibrated granular matter and artificial microswimmers suspended in fluids. The last few decades have seen tremendous progress in understanding their collective properties and phase behaviors [1, 2, 3, 4, 10, 11, 12, 13].
@@
In dry active matter, self-propelled particles (SPP) with local alignment display order-disorder transition coupled with density fluctuations. Consideration of ferromagnetic alignment of the active heading directions of individual propulsion led to flocking and persistent motion of the flocks. This was demonstrated in the celebrated Vicsek model and its variants and described by the Toner-Tu theory of coupled orientation and density fields [14, 15, 16, 17, 4]. Numerical studies of polar rods with volume exclusion showed nematic alignment and clustering [18], and a related simplified model led to system-spanning clusters [19]. A Boltzmann-Landau-Ginzburg kinetic theory approach, based on a Vicsek-like model of active nematics, produced hydrodynamic equations consistent with previous findings [2, 3, 4, 10, 11, 20]. In certain systems, alignment results from physical processes such as actual collisions between active elements [21, 18, 22]. Other instances involve effective interactions mediated by non-equilibrium mechanisms like complex biochemical signaling or visual and cognitive processes [23, 24, 25]. The former alignments are characterized by reciprocal interactions, while non-reciprocal rules often describe the latter [26, 27, 28]. In equilibrium systems, microscopic details are typically considered irrelevant to emergent macroscopic properties if the models share symmetries and conservation laws. However, recent studies on aligning active particles with short-range interactions, be it vectorial or nematic active matter, revealed that microscopic implementations in the form of reciprocity and additivity can influence macroscopic behaviors [23], e.g., leading to first-order or critical order-disorder transition in active apolar nematics in the presence or absence of reciprocity [29].
@@
In this paper, we consider a dry, apolar, active nematics model using a reciprocal additive alignment interaction between local neighbors. An earlier on-lattice implementation using the Lebwohl-Lasher potential of alignment [42] found a fluctuation-dominated phase separation associated with the ordering transition [30]. Our earlier off-lattice calculation of SPPs at a fixed activity demonstrated a clear first-order nematic-isotropic (NI) transition with increasing orientational noise [29].
In this paper, we present a comprehensive phase diagram as a function of changing activity and relative orientational noise with respect to the alignment strength. It shows three different parameter regimes characterized by (i) a homogeneous isotropic phase, (ii) a nematically aligned phase with significant density fluctuations, and (iii) a coexistence of nematic and isotropic in the presence of the largest density fluctuations. The order-disorder transition remains fluctuation-dominated and first-order across all parameter regimes. Using a mean-field hydrodynamic argument, we obtain an analytic prediction for the phase boundary that shows reasonable agreement with our numerical results. Our primary contribution is this direct calculation of the detailed phase diagram.
@@
The rest of the paper is structured as follows: Section 2 details the model of dry active nematic with reciprocal alignment. Section 3 presents the comprehensive phase diagram. Section 4 describes the mean field hydrodynamic analysis, explaining the phase boundaries and the nature of the nematic-isotropic transition. We then explore the coarsening of nematic order and density field in Section 5. Finally, in Section 6, we conclude by summarizing our main results and presenting an outlook.
@@ 2 Model
We adopt periodic boundary conditions. The polarity $q_{i}=\pm 1$ is randomly assigned with equal probability to model apolar particles. The heading direction ${\bf u}_{i}=(\cos\theta_{i},\sin\theta_{i})$ evolves along with the angle $\theta_{i}$ with respect to the $x$-axis. Dynamics are governed by a competition between inter-particle alignment and orientational noise. In active nematics, alignment interactions are designed to make neighboring particles’ heading directions parallel or anti-parallel with equal likelihood.
@@
Employing the Euler-Maruyama scheme, we perform direct numerical simulations by integrating Eq.(1) and (3) using $dt=7\times 10^{-4}\,D_{r}^{-1}$. Care is taken so that $v_{0}dt<r_{c}$ in order to incorporate the influence of interaction properly. We utilize scaled angular diffusion $D_{\theta}:={2D_{r}}/{\mu J}$ and activity $Pe:=v_{0}/D_{r}r_{c}$ as control parameters to construct the phase diagram, delineating associated phases at a fixed dimensionless number density $\rho={N}/{L^{2}}=\rho_{0}$ with $\rho_{0}r_{c}^{2}=0.4$.
The nematic order parameter is $\langle Q_{\alpha\beta}\rangle$ with $Q^{(i)}_{\alpha\beta}=(u_{i\alpha}u_{i\beta}-\frac{1}{d}\delta_{\alpha\beta})$ for $i$-th particle with $\alpha,\beta$ denoting the components. In 2D,
@@
From numerical simulations, we assess the degree of nematic order by utilizing the positive eigenvalue of $\langle Q_{\alpha\beta}\rangle$, the scalar order parameter:
$S=[\langle\cos(2\theta_{i})\rangle^{2}+\langle\sin(2\theta_{i})\rangle^{2}]^{1/2}.$ (4)
@@
Figure 1: (a) Phase diagram in $(Pe,D_{\theta})$ plane with dimensionless activity $Pe=v_{0}/D_{r}r_{c}$ and orientational noise relative to alignment strength $D_{\theta}=2D_{r}/\mu J$ showing three distinct regions: disordered (D), coexistence (C) and nematic (N). The solid black line has the scaling form $D_{\theta}^{*}\sim Pe^{2}$ (see Sec.4 for derivation). (b) The nematic order $S$ (red line points) and standard deviation $\sigma_{s}$ (blue line points) as a function of $D_{\theta}$ at a fixed $Pe=428.57$. (c) The local probability distribution $P(S)$ at $Pe=428.57$ and various $D_{\theta}$ mentioned in the figure legend. The distributions are calculated using $S$ calculated over local coarse-grained volumes of linear dimension $10\,r_{c}$. (e-i) Representative configurations at parameter values marked by solid green boxes in the phase diagram in (a). The parameter values are ${}(Pe,D_{\theta})=$ (142.85, 0.466), (214.28, 0.466), (714.28, 0.254), (714.28, 1.166), (285.71, 1.166), (214.28,1.166), respectively. The color palette to the left of Figure (d) denotes the heading directions of the particles in all the snapshots shown here.
@@ 3 Phase diagram
We begin our investigation by studying the steady states of the model in square geometry of size $L=200$ and $N=16000$, utilizing periodic boundary conditions (PBC). Fig.1(a) shows the phase diagram of the nematic-isotropic (NI) transition. The regions of the nematic phase (N) are denoted by blue diamonds, the region of nematic-isotropic coexistence (C) is indicated by red crosses, and the disordered isotropic phase (D) regions are marked by white open circles. The phase boundary between isotropic and coexistence is denoted by black-filled circles. The heat map in the phase diagram shows the amount of density fluctuations. This is calculated in coarse-grained volumes of linear dimension $10r_{c}$ using $\delta\rho=[\langle\rho^{2}\rangle-\langle\rho\rangle^{2}]^{1/2}$, and then averaging over the system size and all steady state configurations. As the color codes show, the density fluctuation is maximum in the coexistence. Its presence is significant even in the ordered nematic phase. The fluctuation is minimal in the homogeneous isotropic phase.
@@
The three insets show the typical distributions of the scalar nematic order parameter corresponding to nematic (N), coexistence (C), and isotropic (D) regions. An analytic description of the NI phase boundary $D_{\theta}\sim Pe^{2}$ is provided in the following section. Shown by a black solid line in Fig.1(a), it qualitatively captures the actual phase boundary up to $Pe\approx 300$. The broken solid line indicates a breakdown of this simple estimate. Beyond this point, the phase boundary becomes approximately independent of $Pe$. As expected, the active nematic phase is characterized by a giant number fluctuation [29, 30, 2]. For example, deep inside the nematic phase, at $Pe=714.28$ and $D_{\theta}=0.254$, the system shows $\langle\Delta n\rangle^{2}=\langle n^{2}\rangle-\langle n\rangle^{2}\sim\langle n\rangle^{a}$ with $a\approx 1.52$; see Appendix-A. The steady-state pair correlation function $C(r)=1-b(r/L)^{\alpha}$ with $\alpha\approx 0.26$ [29] signifies a clear violation of Porod’s law in the coexistence region. This captures a huge roughness of interfaces evidenced by large $(1-\alpha)$ characterizing a fluctuation-dominated phase separation [44].
@@
The NI transition is captured using the dependence of the scalar order parameter $S$ and its standard deviation $\sigma_{s}=[\langle S^{2}\rangle-\langle S\rangle^{2}]^{1/2}$ on $D_{\theta}$, at a fixed $Pe$; see Fig.1(b). The cross-correlation between fluctuations in nematic order and density peaks near the transition, as detailed in Appendix-B. However, this correlation diminishes significantly within the nematic phase and approximately vanishes after transitioning to isotropic. The local probability distribution $P(S)$ in Fig.1(c) shows an unimodal distribution with the maximum at small $S$ for large $D_{\theta}$, transforms into a bimodal distribution at intermediate $D_{\theta}$ values. This clearly indicates a phase coexistence, characterizing a first-order transition. At an even smaller $D_{\theta}$, the distribution becomes unimodal again, this time displaying a single maximum at large $S$, indicating a nematic phase.
@@
Representative configurations are shown in Fig.1(d)-(i), with orientations distinguished by a color palette shown on the left of Fig.1(d). The parameter values associated with these configurations are marked by solid green squares on the phase diagram Fig.1(a). A typical configuration of the isotropic phase in Fig.1(d) shows an approximately homogeneous distribution of particles and their random orientations of heading direction. At a higher $Pe$, in Fig. 1(e), we observe a local clustering of particles in oriented bands. These bands display a higher nematic order. At an even higher $Pe$ in Fig.1(f), simulations show system-spanning bands with large nematic order coexisting with a uniform isotropic background. This corresponds to the bimodal distribution in $P(S)$ shown in Fig.1(c). Decreasing $D_{\theta}$, keeping $Pe$ unchanged, in Fig.1(g), we observe nice nematic order all through the system, corresponding to a homogeneous nematic fluid phase. The corresponding distribution $P(S)$ is unimodal with the maximum at large $S$. Again, in Fig.1(h), we show a typical configuration at a smaller $Pe$ and $D_{\theta}$, near the phase boundary, to find local nematic bands coexisting with an isotropic background. At an even smaller $Pe$, the system gets back to the isotropic phase in the presence of large density fluctuations; see Fig.1(i).

### v3

# Activity-induced phase transition and coarsening dynamics in dry apolar active nematics

@@ Abstract
Using the Lebwohl-Lasher interaction for reciprocal local alignment, we present a comprehensive phase diagram for a dry, apolar, active nematic system using its stochastic off-lattice dynamics. The nematic-isotropic transition in this system is first-order and occurs alongside a fluctuation-dominated phase separation. Our phase diagram identifies three distinct regions based on activity and orientational noise relative to alignment strength: a homogeneous isotropic phase, a nematic phase with giant density fluctuations, and a coexistence region. Using mean-field analysis and hydrodynamic theory, we demonstrate that reciprocal interactions lead to a density fluctuation-induced first-order transition and derive a phase boundary consistent with numerical results. Quenching from the isotropic to nematic phase reveals coarsening dynamics where nematic ordering precedes particle clustering. Both the nematic and density fields exhibit similar scaling behaviors, exhibiting dynamic exponents $z_{S}\approx 2.5$ and $z_{\rho}\approx 2.34$, consistently falling within the range of 2 and 3.
@@ 1 Introduction
Active nematics represent a fascinating class of nonequilibrium systems characterized by the collective motion and organization of self-propelled entities [1, 2, 3, 4, 5, 6, 7]. Inspired by a myriad of biological systems, ranging from the dynamic organization of the cytoskeleton within cells to the emergent behaviors observed in tissues and even granular materials, active nematics have emerged as a pivotal area of study at the interface of physics, biology, and materials science [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. At the heart of active systems lies the concept of self-propulsion, where individual constituents continually convert internal energy into directed motion. This intrinsic activity, prevalent in biological systems such as motile cells and bacterial colonies, fuels a rich array of dynamic phenomena, including the spontaneous formation of orientational order, collective motion, and the emergence of complex spatiotemporal patterns [4, 8, 9]. Moreover, active nematics find analogs in nonliving systems, including vibrated granular matter and artificial microswimmers suspended in fluids. The last few decades have seen tremendous progress in understanding their collective properties and phase behaviors [1, 2, 3, 4, 7, 16, 19, 17, 20, 21, 22, 23].
@@
In dry active matter, self-propelled particles (SPP) with local alignment display order-disorder transition coupled with density fluctuations. Consideration of ferromagnetic alignment of the active heading directions of individual propulsion led to flocking and persistent motion of the flocks. This was demonstrated in the celebrated Vicsek model and its variants and described by the Toner-Tu theory of coupled orientation and density fields [24, 25, 26, 27, 4]. Self-propelled rods with nematic alignemnt [28, 29, 30, 31, 32, 7] and recent experimental studies [33] display an isotropic-to-nematic transition, with the formation of nematic bands first observed in Vicsek-like models of polar active nematics [34]. Numerical studies of self-propelled rods [29, 34, 35] as well as experiments on M. xanthus in Ref. 36 showed the formation of polar clusters. The overall polar order vanishes in large-scale simulations of related models that led to long-range nematic order in 2D [34]. Recent experiments in actomyosin motility assays [33] indicate that both polar or nematic order can emerge depending on control parameters. A Boltzmann-Landau-Ginzburg kinetic theory approach, based on a Vicsek-like model of active nematics, produced hydrodynamic equations consistent with previous findings [2, 3, 4, 31, 16, 17, 18]. In certain systems, alignment results from physical processes such as actual collisions between active elements [37, 29, 38, 11, 12, 14]. Other instances involve effective interactions mediated by non-equilibrium mechanisms like complex biochemical signaling or visual and cognitive processes [39, 40, 41]. The former alignments are characterized by reciprocal interactions, while non-reciprocal rules often describe the latter [42, 43, 44]. In equilibrium systems, microscopic details are typically considered irrelevant to emergent macroscopic properties if the models share symmetries and conservation laws. However, recent studies on aligning active particles with short-range interactions, be it vectorial or nematic active matter, revealed that microscopic implementations in the form of reciprocity and additivity can influence macroscopic behaviors [39, 45]. For example, active apolar nematics in the presence (or absence) of reciprocity show first-order (or continuous) nematic-isotropic (NI) transition in the presence of fluctuation-dominated phase separation (FDPS) of high and low-density regions [45]. These findings align with the non-reciprocal Vicsek-like models [21, 20], exhibiting an unimodal distribution of nematic order across the transition, indicative of a continuous NI transition, despite in the presence of a coexistence of low-density regions with unstable high-density nematic bands.
@@
In this paper, we consider a dry, apolar, active nematics model using a reciprocal additive alignment interaction between local neighbors. This model is idealized; real systems with active polymers, F-actins, or microtubules may display additional properties, such as emergent elasticity or chiral segregation, which are not covered here [13, 11, 12, 23]. An earlier on-lattice implementation of active Lebwohl-Lasher model [56] found an FDPS associated with the ordering transition [46]. Our earlier off-lattice calculation of SPPs at a fixed activity demonstrated a clear first-order NI transition with increasing orientational noise [45].
In this paper, we present a comprehensive phase diagram as a function of changing activity and relative orientational noise with respect to the alignment strength. Previous studies on SPP often used density and noise as control parameters [17, 20]. This work emphasizes the recurring theme of order emerging from activity [28, 30], demonstrating how the strength of activity influences the phase diagram. It shows three different parameter regimes characterized by (i) a homogeneous isotropic phase, (ii) a nematically aligned phase with significant density fluctuations, and (iii) a coexistence of nematic and isotropic in the presence of the largest density fluctuations. The phase transition is characterized by FDPS and first-order NI transition across all parameter regimes. Using a mean-field hydrodynamic argument, we obtain an analytic prediction for the phase boundary that shows reasonable agreement with our numerical results. Our primary contribution is this direct calculation of the detailed phase diagram.
@@
The rest of the paper is structured as follows: Section 2 details the model of dry active nematic with reciprocal alignment. Section 3 presents the comprehensive phase diagram. Section 4 describes the mean field hydrodynamic analysis, explaining the phase boundaries and the nature of the NI transition. We then explore the coarsening of nematic order and density field in Section 5. Finally, in Section 6, we conclude by summarizing our main results and presenting an outlook.
@@ 2 Model
We adopt periodic boundary conditions. The polarity $q_{i}=\pm 1$ is randomly assigned with equal probability, at each time step to model apolar particles. The heading direction ${\bf u}_{i}=(\cos\theta_{i},\sin\theta_{i})$ evolves along with the angle $\theta_{i}$ with respect to the $x$-axis. Dynamics are governed by a competition between inter-particle alignment and orientational noise. In active nematics, alignment interactions are designed to make neighboring particles’ heading directions parallel or anti-parallel with equal likelihood.
@@
Employing the Euler-Maruyama scheme, we perform direct numerical simulations by integrating Eq.(1) and (3) using $dt=7\times 10^{-4}\,D_{r}^{-1}$. Care is taken so that $v_{0}dt<r_{c}$ to incorporate the influence of interaction properly ; the condition ensures that interaction between particles is well resolved in the dynamics [68]. We utilize the scaled angular diffusion $D_{\theta}:={2D_{r}}/{\mu J}$ and activity in terms of Péclet number $Pe:=v_{0}/D_{r}r_{c}$ as dimension-less control parameters to construct the phase diagram, delineating associated phases at a fixed dimensionless number density $\rho={N}/{L^{2}}=\rho_{0}$ with $\rho_{0}r_{c}^{2}=0.4$. The scaled angular diffusion $D_{\theta}$ is a ratio of orientational relaxation time $(\mu J)^{-1}$ due to alignment interaction to the single-particle persistence time $D_{r}^{-1}$. The Péclet number $Pe$ is defined as a ratio of active persistence length $v_{0}/D_{r}$ of a single SPP to the alignment interaction range $r_{c}$, to capture the effect of interaction.
[…]
@@
From numerical simulations, we determine the degree of nematic order by analyzing the positive eigenvalue of $\langle Q_{\alpha\beta}\rangle$, which represents the scalar order parameter  (see Appendix-A for further details) :
$S=\left[\langle\cos(2\theta_{i})\rangle^{2}+\langle\sin(2\theta_{i})\rangle^{2}\right]^{1/2}.$ (4)
@@
Figure 1: (a) Phase diagram in $(Pe,D_{\theta})$ plane with dimensionless activity $Pe=v_{0}/D_{r}r_{c}$ and orientational noise relative to alignment strength $D_{\theta}=2D_{r}/\mu J$ showing three distinct regions: isotropic homogeneous (Isotropic), nematic-isotropic coexistence (Coexistence) and nematically ordered (Nematic). The solid black line has the scaling form $D_{\theta}^{*}\sim Pe^{2}$ (see Sec.4 for derivation). (b) The nematic order $S$ (red line points) and standard deviation $\sigma_{s}$ (blue line points) as a function of $D_{\theta}$ at a fixed $Pe=428.57$. (c) The local probability distribution $P(S)$ at $Pe=428.57$ and various $D_{\theta}$ mentioned in the figure legend. The distributions are calculated using $S$ calculated over local coarse-grained volumes of linear dimension $10\,r_{c}$. (d-i)  Representative configurations at parameter values marked by solid green boxes in the phase diagram in (a). The parameter values are ${}(Pe,D_{\theta})=$ (214.28, 1.166), (285.71, 1.166), (714.28, 1.166), (714.28, 0.254), (214.28, 0.466), (142.85, 0.466) respectively. The color palette to the left of Figure (d) denotes the heading directions of the particles in all the snapshots shown here.
@@ 3 Phase diagram
We begin our investigation by studying the steady states of the model in square geometry of size $L=200$ and $N=16000$, utilizing periodic boundary conditions (PBC). Fig.1(a) shows the phase diagram of the NI transition. The regions of the nematic phase ( Nematic ) are denoted by blue diamonds, the region of NI coexistence ( Coexistence ) is indicated by red crosses, and the homogeneous isotropic phase ( Isotropic ) region is marked by white open circles. The phase boundary between Isotropic and Coexistence is denoted by black-filled circles. The phase diagram clearly demonstrates that nematic order arises from activity: the system remains in an isotropic phase at $Pe=0$, and only when $Pe$ exceeds a certain threshold does nematic order begin to emerge. The heat map in the phase diagram shows the amount of density fluctuations. This is calculated in coarse-grained volumes of linear dimension $10r_{c}$ using $\delta\rho=[\langle\rho^{2}\rangle-\langle\rho\rangle^{2}]^{1/2}$, and then averaging over the system size and all steady state configurations. As the color codes show, the density fluctuation is maximum in the coexistence. Its presence is significant even in the ordered nematic phase. The fluctuation is minimal in the homogeneous isotropic phase.
@@
The three insets in Fig.1(c) illustrate the typical distributions of the scalar nematic order parameter, corresponding to nematic, coexistence, and isotropic regions. An analytic description of the NI phase boundary $D_{\theta}\sim Pe^{2}$ is provided in the following section. Shown by a black solid line in Fig.1(a), it qualitatively captures the actual phase boundary up to $Pe\approx 300$. The broken solid line indicates a breakdown of this simple estimate. Beyond this point, the phase boundary becomes approximately independent of $Pe$. Unlike in equilibrium systems, where number fluctuations follow $\langle\Delta n^{2}\rangle=\langle n^{2}\rangle-\langle n\rangle^{2}\sim\langle n\rangle^{a}$ with $a=1$, active nematic systems exhibit $a>1$, known as giant number fluctuations [1, 45, 46]. Consistent with this expectation, we observe $a\approx 1.52$ in the nematic phase at $Pe=714.28$ and $D_{\theta}=0.254$; details can be found in Appendix-B. Additionally, as previously demonstrated [45], the active nematic model employing the Lebwohl-Lasher potential exhibits FDPS [58].
@@
The NI transition is captured using the dependence of the scalar order parameter $S$ and its standard deviation $\sigma_{s}=[\langle S^{2}\rangle-\langle S\rangle^{2}]^{1/2}$ on $D_{\theta}$, at a fixed $Pe$; see Fig.1(b). The cross-correlation between fluctuations in nematic order and density peaks near the transition, as detailed in Appendix-C. However, this correlation diminishes significantly within the nematic phase and approximately vanishes after transitioning to isotropic. The local probability distribution $P(S)$ in Fig.1(c) shows an unimodal distribution with the maximum at small $S$ for large $D_{\theta}$, transforms into a bimodal distribution at intermediate $D_{\theta}$ values. This clearly indicates a phase coexistence, characterizing a first-order transition. At an even smaller $D_{\theta}$, the distribution becomes unimodal again, this time displaying a single maximum at large $S$, indicating a nematic phase.
@@
Representative configurations are shown in Fig.1(d)-(i), with orientations distinguished by a color palette shown on the left of Fig.1(d). The parameter values associated with these configurations are marked by solid green squares on the phase diagram Fig.1(a). A typical configuration of the isotropic phase in Fig.1(d) shows an approximately homogeneous distribution of particles and their random orientations of heading direction. At a higher $Pe$, in Fig. 1(e), we observe a local clustering of particles in oriented bands. These bands display a higher nematic order. At an even higher $Pe$ in Fig.1(f), simulations show system-spanning bands with large nematic order coexisting with a uniform isotropic background. This corresponds to the bimodal distribution in $P(S)$ shown in Fig.1(c). However, the nematic bands show transverse instability; they form and disappear [59, 46, 19, 17, 32]. Appendix F presents configurations as bands form and break. Decreasing $D_{\theta}$, keeping $Pe$ unchanged, in Fig.1(g), we observe nice nematic order all through the system, corresponding to a homogeneous nematic fluid phase. The corresponding distribution $P(S)$ is unimodal with the maximum at large $S$. Again, in Fig.1(h), we show a typical configuration at a smaller $Pe$ and $D_{\theta}$, near the phase boundary, to find local nematic bands coexisting with an isotropic background. At an even smaller $Pe$, the system gets back to the isotropic phase in the presence of large density fluctuations; see Fig.1(i).